A ball is dropped from a high rise platform at t=0 starting from rest . After 6 seconds another ball is thrown downwards from the same platform with a speed v . The two balls meet at t =18 s . What is the value of v.

Grade:11

5 Answers

Hasan Naqvi
8 years ago
Distance between the two balls after 6 secs = ½ * 10 * 62 = 180
Relative velocity of second ball(wrt first ball) so that it can cover this distance = 180/ 12 = 15m/s
 
Velocity of first ball (at t= 6) =  10* 6 = 60
 
v= 60 + 15 =75m/s
Soumyadeep
8 years ago
How come it's 180/12 
Nikita Agarwal
6 years ago
The distance travelled by 1st ball in 18 sec is h=1/2gt^2=1/2*10*18^2=1620mNow to meet 2nd ball has to cover the same distance in (18-6 ) sec =12secSo for 2nd ballh=vt+1/2gt^21620=v*12+1/2*10*12^2=75m/s
khushi jangir
6 years ago
diatance travelled by 1st ball in 18 sec=ut+1/2at^2s=0+1/2×10×18×18s=90×18s=1620mtime taken by second ball=18-6=12secso,distance travelled by second ball=ut+1/2gt×tso,1620==12×u+1/2×10×12×12u=135-60u=75m
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
since both cover the same distance downwards till t = 18
hence, h = ½ g (18)2 = vt + ½ g t2
where, t = 18 – 6 = 12s
hence, v * 12 + ½ g(12)2 = ½ g 182
v = ½ g (182 – 122)/12
v = 0.5 * 9.81 * (324 – 144)/ 12
v = 73.575 m/s
 
Hope it helps.
Thanks and regards,
Kushagra